OpenStudy (ineedhelp10):

Two cars travel the same distance. Car 1 travels at 50 mph and Car 2 at 65 mph. If it takes Car 1 one more hour to travel than Car 2, how far did the cars travel? Remember d=rt.

OpenStudy (isaidavila):

Do you have choices?

OpenStudy (3redsons):

is it how far they travel together?

OpenStudy (3redsons):

LIKE THE TOTAL DISTANCE sorry caps

OpenStudy (isaidavila):

That's what I get D=D V1*T1=V2*T2 50(t+1)=65t 50t+50=65t 50=15t t=10/3

OpenStudy (3redsons):

i do not have enough information

OpenStudy (ineedhelp10):

It's not multiple choice but I have to answer different parts, which are. A. Write an expression to represent the distance car 1 and car 2 travel B. Use these expressions to find Car 2 traveled

OpenStudy (ineedhelp10):

The final question is to find the distance traveled

OpenStudy (3redsons):

i know that car 1 needs to be one more hour, but what was the original hour

OpenStudy (isaidavila):

Since both cat travelled the same distance D=65.10/3=130

OpenStudy (ineedhelp10):

It doesn't say the original hour all it says is the speed they were both going

OpenStudy (3redsons):

ok 2 hard 4 meee, sorry

OpenStudy (ineedhelp10):

Help